If the force of attraction between two unit masses (1 kg each) kept 1 km apart is taken as unit force and the unit of energy in this system is called ‘eluoj’ (joule written in reverse order) then the relation of ‘joule’ and ‘eluoj’ is [G = 6.67 × 10–11 Nm2 kg]–2


        (a)    1 eluoj = 6.67 × 10–17 joule


        (b)    1 eluoj = 6.67 × 10–14 joule


        (c)    1 eluoj = 10–4 joule


        (d)    1 eluoj = 6.67 × 10–9 joule


Grade:upto college level

2 Answers

Simran Bhatia
10 years ago

(b)

 

        Let’s say force ‘notwen’ (newton written in reverse order) we know

 

        1 notwen = G  = G = 106 notwen m2kg–2

 

        But G = 6.67 × 10–11 Nm2kg–2

 

        \ 1 notwen = 6.67 × 10–17 newton

 

        1 eluoj = 1 Nowten × 1 km = 6.67 × 10–14 newton × 1 m

 

= 6.67 × 10–14 Joule

Arnab Mukherjee
6 years ago
(b) Let’s say force ‘notwen’ (newton written in reverse order) we know 1 notwen = G = G = 106 notwen m2kg–2 But G = 6.67 × 10–11 Nm2kg–2 \ 1 notwen = 6.67 × 10–17 newton 1 eluoj = 1 Nowten × 1 km = 6.67 × 10–14 newton × 1 m = 6.67 × 10–14 Joule